# Observational constraints on generalised axion-like potentials for the late Universe

### v1

@@ 4 Conclusions
We have fitted the axion-inspired scalar field dark energy model with a potential of the inverse power of cosine in Eq. (2.3) from [36] using the complete cosmological datasets of CMB (Planck2018 $-\ell$ TTEE [39], NPIPE CamSpec $-\ell$ TTTEEE [16], PR4 CMB lensing [40, 41]) , BAO (DESI DR1 without full-shape [42, 43, 44]), supernovae (Pantheon+ SNe catalog [18] without anchoring of SNe standardised absolute magnitude), low-$z$ $H_{0}$ measurements (Riess et al. low-$z$ anchors [17]), and DES Y1 morphology [45]. In addition to the existence of the tracking solution common to many scalar field dark energy models that solves the initial condition problem, the model considered in this work aims to solve the coincidence problem of why the dark energy only dominates at $z\lesssim 1$ with $w\sim-1$ by offering a graceful exit of the tracking regime. The model seems promising from the perspective of naturalness in parameter selection, as the system closely follows the $\Lambda$CDM model for the majority of the time, except during the matter dominated era. We hoped that the model may ease the $S_{8}$ tension without affecting the Hubble tension through an early matter-dark energy transition, as demonstrated in [37].
Unfortunately, the model shows no statistically meaningful benefit over $\Lambda$CDM model across Planck, DESI, Pantheon+, Riess et al. low-$z$ anchors and DES Y1 morphological datasets. A reduction in tension by $0.1\pm 0.3\sigma$ is realised between high-$z$ dataset of Planck and DESI, and low-$z$ dataset of Pantheon+ and $H_{0}$ anchors, at the cost of an increase in tension by $0.3\pm 0.5\sigma$ between the combined dataset of Planck, DESI, Pantheon+, $H_{0}$ anchors, and the low-$z$ matter perturbation dataset of DES Y1. This clearly cannot solve the cosmological tension. In addition, the structure formation history remains practically identical to the one in $\Lambda$CDM model. We isolate the root cause to the inability of the axion-like dark energy model to increase $H_{0}$, and thus decrease $\Omega_{m0}$ under the constraint of Planck CMB data. The salient feature of the model, a long-lasting $w>-1$ tracking regime during the matter-dominated era with a graceful exit, is thus rendered ineffective.

### v2

@@ 4 Conclusions
We have fitted the axion-inspired scalar field dark energy model with a potential of the inverse power of cosine in Eq. (2.3) from [36] using the complete cosmological datasets of CMB (Planck2018 $-\ell$ TTEE [39], NPIPE CamSpec $-\ell$ TTTEEE [16], PR4 CMB lensing [40, 41]) , BAO (DESI DR1 without full-shape [42, 43, 44]), supernovae (Pantheon+ SNe catalog [18] without anchoring of SNe standardised absolute magnitude), low-$z$ $H_{0}$ measurements (Riess et al. low-$z$ anchors [17]), and DES Y1 morphology [45]. In addition to the existence of the tracking solution common to many scalar field dark energy models that solves the initial condition problem, the model considered in this work aims to solve the coincidence problem of why the dark energy only dominates at $z\lesssim 1$ with $w\sim-1$ by offering a graceful exit of the tracking regime. The main attraction of the model is its insensitivity to the model parameters, as the system closely follows the $\Lambda$CDM model, except for a finite duration after matter-radiation crossover. We hoped that the model may ease the $S_{8}$ tension without affecting the Hubble tension through an early matter-dark energy transition, as demonstrated in [37].
The model shows no statistical disadvantage nor meaningful advantage over $\Lambda$CDM model across datasets considered in this work. A reduction in tension by $0.1\pm 0.3\sigma$ is realised between high-$z$ dataset of Planck and DESI, and low-$z$ dataset of Pantheon+ and $H_{0}$ anchors, but is compensated by an increase of $0.3\pm 0.5\sigma$ in the tension with the low-$z$ weak lensing dataset of DES Y1. In particular, the structure formation history remains practically identical to the one in $\Lambda$CDM model. We isolate the root cause to the inability of the axion-like dark energy model to increase $H_{0}$, and thus decrease $\Omega_{m0}$ under the constraint of Planck CMB data. The salient feature of the model, a long-lasting $w>-1$ tracking regime during the matter-dominated era with a graceful exit, is thus rendered ineffective.
@@ Appendix A Statistical probes
Appendix B Prior volume bias analysis
$\phi$CDM $\Lambda$CDM
Parameters CMB + BAO + SNe CMB + BAO +SNe
$\log(10^{9}A_{s})$ $-0.16$ $+0.11$ $-0.05$ $+0.19$ $-0.04$ $-0.32$
$n_{s}$ $+0.39$ $+0.20$ $-0.19$ $-0.40$ $+0.22$ $+0.27$
$\Omega_{b0}h^{2}$ $+0.31$ $+0.34$ $+0.06$ $-0.03$ $+0.15$ $-0.55$
$\Omega_{c0}h^{2}$ $-0.70$ $-0.11$ $+0.06$ $+0.15$$\pm 0.00$ $-0.12$
$\tau_{\mathrm{reio}}$ $-0.08$ $+0.09$ $-0.15$ $+0.03$ $+0.09$ $-0.25$
$100\theta_{MC}$ $+0.12$ $-0.33$ $+0.75$ $+0.35$ $+0.75$ $-0.13$
$y_{\mathrm{cal}}$ $+0.06$ $+0.31$ $+0.02$ $+0.26$ $-0.38$ $-0.05$
$A_{143}$ $-0.31$ $+0.32$ $-0.13$ $+0.32$ $-0.03$ $+0.22$
$A_{217}$ $-0.00$ $+0.15$ $-0.29$ $+0.25$ $+0.06$ $+0.25$
$A_{143\times 217}$ $-0.13$ $+0.21$ $-0.23$ $+0.24$ $-0.04$ $+0.17$
$\gamma_{143}$ $+0.33$ $+0.04$ $+0.22$ $-0.08$ $-0.14$ $-0.12$
$\gamma_{217}$ $-0.05$ $-0.21$ $+0.31$ $-0.07$ $-0.15$ $-0.35$
$\gamma_{143\times 217}$ $+0.02$ $-0.34$ $+0.22$ $-0.38$ $-0.36$ $-0.44$
$c_{TE}$ $-0.37$ $+0.25$ $+0.20$ $+0.17$ $+0.24$ $+0.61$
$c_{EE}$ $-0.27$ $-0.24$ $+0.18$ $+0.69$ $-0.03$ $+0.15$
$H_{0}$ $+0.74$ $+0.32$ $+0.15$ $-0.07$ $+0.20$ $-0.09$
$\Omega_{m0}$ $-0.76$ $-0.26$ $-0.11$ $+0.09$ $-0.14$ $-0.02$
$S_{8}$ $-0.81$ $-0.04$ $-0.04$ $+0.18$ $-0.08$ $-0.19$
$\mathcal{E}_{H_{0}}$ $+0.03$ $-0.58$ $-1.07$
$n$ $-0.69$ $+0.79$ $-0.87$
$\log_{10}(\rho_{\mathrm{DE,0}}/V_{\mathrm{min}}-1)$ $-0.89$ $-1.54$ $+1.29$
$\log_{10}(n\eta^{-2})$ $+0.95$ $+0.86$ $-1.01$
$\log_{10}(\phi_{i}/\eta)$ $+0.79$ $-0.10$ $+0.58$
Table 8: The shift of model and nuisance parameters between maximum-a-posteriori (MAP) value and Bayesian mean in unit of Bayesian standard deviation $(\theta_{\mathrm{map}}-\left\langle\theta\right\rangle)/\sqrt{\left\langle\theta^{2}\right\rangle-\left\langle\theta\right\rangle^{2}}$ for the axion-like dark energy model ($\phi$CDM) and $\Lambda$CDM model in CMB, CMB + BAO, and CMB + BAO + SNe dataset. Deviation between $0.45\sim 0.65\sigma$ and above is marked orange and red respectively.
As made apparent in Figs. 6 and 7, the constraint on the extended parameters is rather weak. This makes the choice of the prior of these parameters in table 2 the dominant factor in Bayesian analysis. Different priors, especially those biasing toward specific direction, could lead to Bayesian mean values different from the maximal a posteriori (MAP) value, signalling the breakdown of the Gaussian approximation. We verify that the Bayesian mean does not deviate far away from MAP value in the analysis of section 3 as shown in table 8 other than the extended parameters and the Gaussian constraint $\mathcal{E}_{H_{0}}$, with the sole exception of the axion-like dark energy model against the CMB dataset where the distribution of multiple observables does become non-Gaussian.
This is consistent with the conclusion drawn from Fig. 7 in section 3.2, that without any late-time measurement, the CMB dataset alone permits strong modification of the cosmic evolution, i.e., a dark energy with an equation of state deviating away from $-1$ during the dark-energy-dominated era. The finite prior volume of $n$ thus biases the posterior toward $\Lambda$CDM-like configurations ($n\sim 0$). This bias is further enhanced by the finite width of the Gaussian constraint that greatly expands the prior volume around $\log_{10}(\rho_{\rm DE,0}/V_{\rm min}-1)\sim-3$ and $n\eta^{-2}\gg 1$.
We emphasise again that this biasing effect on the Bayesian analysis is fully anticipated, and is in fact the key feature of the axion-like dark energy model considered in this work, i.e. a mechanism for the grand entrance and graceful exit of the tracking regime. The sliding of the tracking regime guarantees a finite window within an infinite parameter space where the axion-like dark energy behaves differently than a cosmological constant. Therefore, the prior volume ratio between that finite window and the infinite $\Lambda$CDM-like region will always introduce bias into the analysis.
Figure 12: Scatter plot of the late-time parameters, the extended parameters and the Gaussian constraint for the axion-like dark energy model against CMB + BAO + SNe dataset, colour-coded by $\log_{10}(\rho_{\rm DE,0}/V_{\rm min}-1)$.
Figure 13: Scatter plot of the late-time parameters, the extended parameters and the Gaussian constraint for the axion-like dark energy model against CMB + BAO + SNe + low-$z$ dataset, colour-coded by $\log_{10}(\rho_{\rm DE,0}/V_{\rm min}-1)$.
As evidently shown in table 8, such a bias toward $\Lambda$CDM is clearly suppressed once considering distance anchors, as the dark energy equation of state is constrained to be close to $-1$. However, the finite width effect of the Gaussian constraint remains, as it introduces bias toward $n\eta^{-2}\gg 1$ as show in Figs. 12 and 13. This again leads to bias toward $\Lambda$CDM in the case of CMB + BAO + SNe dataset, but not in the case of CMB + BAO + SNe + low-$z$ dataset as the distance measurements are finally strong enough to exclude all configurations of the axion-like dark energy that behaves differently from $\Lambda$CDM in the dark-energy-dominated era.